Write four functions: (IN PYTHON 3)

1) bound(l) - given a list of integers l, compute a tuple containing the minimum and maximum values (in that order)

2) span(l) - given a list of integers l, compute the difference between the minimum and maximum function.

3) mean(l) - given a list of integers l, compute the average. The return value will be a float.

4) stats(f, l) - given a function f and a list l, apply the function f to the list and return the result.
Answer #1
  1. bound(l) function

This receives a list l as parameter, and finds the minimum and maximum of the list and return the tuple of it.

This can be written as:

# bound(l) - given a list of integers l, compute a tuple containing the minimum and maximum values (in that order)

def bound(l):

    # get minimum of the list

    minimum = min(l)

    # get maximum of the list

    maximum = max(l)

    # return the tuple containing the minimum and maximum values

    return (minimum, maximum)

Below is the sample output:

  1. span(l) function

This function receives a list and find the minimum and maximum of the list like the above function. Then it calculates the difference between maximum and minimum and return that value.

This could be coded as:

# span(l) - given a list of integers l, compute the difference between the minimum and maximum function.

def span(l):

    # get minimum of the list

    minimum = min(l)

    # get maximum of the list

    maximum = max(l)

    # calculate the difference between maximum and minimum

    diff = maximum - minimum

    # return the difference

    return diff

Below is the sample output:

  1. mean(l) function

This function receives a list and calculate the mean of numbers of list and returns it. To calculate the mean, we iterate through the list and calculate the total of all value, then divide that value by the length of list, which will give the mean and then it returns it.

This could be coded as:

# mean(l) - given a list of integers l, compute the average. The return value will be a float

def mean(l):

    # calculate the total of items

    total = 0

    for item in l:

        total += item

    # get size of list

   size = len(l)

    # calculate the average by dividing the total by size of list

    average = float(total/size)

    # return the average

    return average

Below is the sample output:

  1. stats(f, l) function

This function receives another function and a list as a parameter. It calls the given function with list as its parameter. It returns the return value returned by the function f. It basically can be used to call above functions as its parameter.

This could be coded as:

# given a function f and a list l, apply the function f to the list and return the result

def stats(f, l):

    # get value out of function

    value = f(l)

    # return the value which the return value of the respective function

    return value

Below is the sample output:

All the required functions have been coded and explained above. Let me know if you have any queries.

Thanks!